Explain nominal and real interest rate.What is the Fisher Effect? If=1.7%and=0.8%, computethereal interest rateusing the exact formula and the approximationformula.
1
Expert's answer
2020-05-12T10:45:57-0400

Nominal interest rate refers to the interest rate before taking inflation into account. Unlike the nominal rate, the real interest rate takes the inflation rate into account. 

Fisher Effect states that the real interest rate equals the nominal interest rate minus the expected inflation rate.

If nominal interest rate is i = 1.7% and inflation is pi = 0.8%, then the real interest rate is:

"r = i - pi = 1.7 - 0.8 = 0.9" %.